OpenStudy (anonymous):

Ernesto rode on a train that traveled at a constant rate. The graph compares the number of hours traveled to the number of miles traveled. What was the rate of speed of the train? A. 35 miles per hour B. 70 miles per hour C. 105 miles per hour D. 140 miles per hour

OpenStudy (redplanet2001):

we need to see the graph

OpenStudy (anonymous):

here the graph

OpenStudy (xguardians):

No permission to view that image.

OpenStudy (anonymous):

why

OpenStudy (anonymous):

maybe just use the snipping tool to get the image

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that the best way i can it on here

OpenStudy (anonymous):

much easier:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pick two points from the graph and use the formula for the slope

OpenStudy (anonymous):

can you tell me the formula because i lost the formula

OpenStudy (anonymous):

\(\Large v=\frac{d_2-d_1}{t_2-t_1}=\frac{\Delta d}{\Delta t}\)